# Toward General-Purpose Robots via Foundation Models: A Survey and Meta-Analysis

### v2

@@ 3 Challenges in Robotics
In this section, we summarize five core challenges that various modules in a typical robotic system face, each detailed in the following subsections. Whereas similar challenges have been discussed in prior literature (Section 1.2), this section mainly focuses on the challenges that may potentially be solved by appropriately leveraging foundation models, given the evidence from current research results. We also depict the taxonomy in the section for easier review in Figure 3.
ptChallenges in Robotics (3)ptGeneralization (3.1)ptperceptionptmorphologyptData Scarcity (3.2)ptsimulationptreal-worldptRequirements of Models (3.3)ptmodel-basedptmodel-freeptTask Specification (3.4) ptgoalimageptrewardgenerationptlanguagepromptptUncertainty and Safety (3.5) ptepistemicptaleatoricptprovablesafety
@@ 3.1 Generalization
Robotics systems often struggle with accurate perception and understanding of their environment. Limitations in computer vision, object recognition, and semantic understanding made it difficult for robots to effectively interact with their surroundings. Traditional robotics systems often relied on analytic hand-crafted algorithms, making it challenging to adapt to new or unseen situations. They also lacked the ability to generalize their training from one task to another, further limiting their usefulness in real-world applications. This generalization ability is also reflected in terms of the generalization of planning and control in different tasks, environments, and robot morphologies. For example, specific hyperparameters for, e.g., classical motion planners and controllers need to be tuned for specific environments [164, 102, 103]; RL-based controllers are difficult to transfer across different tasks and environments [47, 165]. In addition, due to differences in robotic hardware, it is also challenging to transfer models across different robot morphologies [166, 167]. By applying foundation models in robotics, the generalization problem is partially resolved, which will be discussed in the next Section 4. Further challenges, such as generalization across different robotic morphologies, remain demanding.
@@ 3.2 Data Scarcity
Data has always been the cornerstone of learning-based robotics methods. The need for large-scale, high-quality data is essential to develop reliable robotic models. Several endeavors have been attempted to collect large-scale datasets in the real world, including autonomous driving [1, 2, 168], robot manipulation trajectories [112, 111, 169], etc. Collecting robot data from human demonstration is expensive [27]. The diverse range of tasks and environments where robots are used even complicates the process of collecting adequate and extensive data in the real world. Moreover, gathering data in real-world settings can be problematic due to safety concerns [164]. To overcome these challenges, many works [170, 171, 172, 173, 174, 175] attempt generating synthetic data in simulated environments. These simulations offer realistic virtual worlds where robots can learn and apply their skills to real-life scenarios. Simulations also allow for domain randomization, as well as the potential to update the parameters of the simulator to better match the real world physics [164], helping robots to develop versatile policies. However, these simulated environments still have their limits, particularly in the diversity of objects, making it difficult to apply the learned skills directly to real-world situations. Collecting real-world robotic data with a scale comparable to the internet-scale image/text data used to train foundation models is especially challenging. One promising approach is collaborative data collection across different laboratories and robot types [176], as shown in Figure 3(a). However, an in-depth analysis of the Open-X Embodiment Dataset reveals certain limitations regarding data type availability. Primarily, the robot morphology utilized for data collection is restrictive; out of the top 35 datasets, 30 are dedicated to single-arm manipulation tasks. Only one dataset pertains to quadruped locomotion, and a single dataset addresses bi-manual tasks. Secondly, the predominant scene type for these manipulation tasks is tabletop setups, often employing toy kitchen objects. Such objects come with inherent assumptions including rigidity and negligible weight, which may not accurately represent a wider range of real-world scenarios. Thirdly, our examination of data collection methods indicates a predominance of human expert involvement, predominantly through virtual reality (VR) or haptic devices. This reliance on human expertise highlights the challenges in acquiring high-quality data and suggests that significant human supervision is required. For instance, the RT-1 Robot Action dataset necessitated a collection period of 17 months, underscoring the extensive effort and time commitment needed for data accumulation with human involvement.
(a) Joint Effort in Real-World Data Collection
(b) Dataset Morphology Distribution
(c) Scene Type Classification
(d) Data Collection Methods
@@
Figure 4: Comprehensive visualizations of the Open-X Embodiment Dataset encompassing data collection methods, robot morphologies, and scene types.
@@ 3.3 Requirements of Models and Primitives
Classical planning and control methods usually require carefully engineered models of the environment and the robot. Optimal control methods require good dynamics models (i.e., world transition models) [8, 177]; motion planning requires a map of the environment [178], the states of the objects robots interact with [179], or a set of pre-defined motion primitives [180]; task planning requires pre-computed object classes and pre-defined rules [92], etc. Previous learning-based methods (e.g., imitation and reinforcement learning) train policies in an end-to-end manner that directly gets control outputs from sensory inputs [112], avoiding building and using models. These methods partially solve this problem of relying on explicit models, but they often struggle to generalize across different environments and tasks. This raises two problems: (1) How can we learn model-agnostic policies that can generalize well? Or, alternatively, (2) How can we learn good world models so that we can apply classical model-based approaches? We see some recent works that aim to resolve these problems using foundation models (especially in a model-free manner), which will be systematically discussed in Section 4. However, the call for world models for robotics remains an intriguing frontier, which will be discussed in Section 6.
@@ 3.4 Task Specifications
Understanding the task specification and grounding it in the robot’s current understanding of the world is a critical challenge to obtaining generalist agents. Often, these task specifications are provided by users with limited understanding of the limitations on the robot’s cognitive and physical capabilities. This not only raises questions about what the best practices are for providing these task specifications, but also about the naturalness and ease of crafting these specifications. Understanding and resolving ambiguity in task specifications, conditioned on the robot’s understanding of its own capabilities, is also challenging. Foundation models, again, are a promising solution for this challenge: task specification can be formulated as language prompts [27, 28, 24], goal images [181], rewards for policy learning [26, 182], etc.
@@ 4 Review of Current Research Methodologies
In this section, we summarize the current research methodologies of foundation models for Robotics. In Section 4.1, we mainly discuss how foundation models are used in robotics in two categories: Foundation Models used in Robotics and Robotics Foundation Models (RFMs). For Foundation Models used in Robotics, we mainly highlight applications of vision and language foundation models used in a zero-shot manner, meaning no additional fine-tuning or training is conducted. In Section 4.2, however, we mainly focus on Robotics Foundation Models, wherein these approaches may warm-start models with vision-language pre-trained initialization and/or directly train the models on robotics datasets. Figure 5 shows the detailed taxonomy of this section.
As introduced in the Section 2 (Preliminaries), a typical robotic system consists of perception, planning, and control modules. In this section, we review the methods presented in these papers following this classification method. Here, we combine motion planning and control into one piece—action generation and treat motion planning modules as higher-level and control as lower-level action generation. It is important to notice that although most of the works use foundation models in different functionality modules of the robotic systems, we will classify these papers into categories based on the module to which the paper contributes the most. There are, however, certain applications of the vision and language foundation models that go across these robotics modules, e.g., grounding of these models in robotics, and generating data from LLMs and VLMs. Given the autoregressive nature of current LLMs, they often grapple with extended horizon tasks. Thus, we also delve into advanced prompting methods proposed in the literature to ameliorate this limitation and enhance planning power. We list these applications in sections 4.1.4, 4.1.5 and 4.1.6, as a different perspective to analyze these works.
@@
ptFoundation Models for Robotics (4) ptRobotic Foundation Models (4.2)ptGeneral-purpose (4.2.2) ptSingle-purpose ptAction Generation Models(4.2.1)ptVision and LanguagePre-training ptReinforcement Learningat Scale pt(Languange-Conditioned)Imitation LearningptMotion Planning Models(4.2.1)pt(Vision Languange)Foundation Modelsused in Robotics (4.1)ptPrompting (4.1.6) ptData Generation (4.1.5) ptGrounding (4.1.4) ptRobotic ModulesptLLM and VLM inAction Generation (4.1.3) ptLLM and VLMin Task Planning (4.1.2) ptVFM and VLMin Perception (4.1.1)
@@ 4.1 Foundation Models used in Robotics
In this section, we focus on zero-shot application of vision and language foundation models in robotics. This mainly includes zero-shot deployment of VLMs used in robotic perception, in context learning of LLMs for task-level and motion-level planning, as well as action-generation. We show a few representative works in Figure 5(c).
@@ 4.1.1 VFMs and VLMs in Robot Perception
Recently, the grounding of vision and language foundation models with geometric and object-centric representations of the world has enabled tremendous progress in context understanding, which is a vital requirement for robots to interact with the real world. We will thoroughly examine the application of VFMs and VLMs in robotic perception from various perspectives.
(a) Object recognition: CLIPort [196]
(b) Object recognition & localization: NLMap [25]
(c) Feature distillation in NeRF: F3RM [197]
Figure 6: Using VLMs and VFMs for robotic perception, in various applications. Due to the limit of space and for the consistency of the figures, we only list part of the works we discussed in this figure.
VLM used for: Object and Scene Representation VFM and VLM used for: Policy learning, State estimation and Interactive perception
(d) Representation in policy learning (also applied feature distillation in Ge-NeRF): GNFactor [198]
(e) State estimation: FM-Loc [199]
(f) Interactive perception: MOSAIC [200]
@@
Most of the previous works discussed above utilize only 2D representation of the objects and environment. To enrich the representation of foundation models in 3D space, F3RM [197] and GNFactor [198] distill 2D foundation model features into 3D space, by combining with NeRF and generalizable NeRF. In addition, GNFactor [198] also apply these distilled features in policy learning. Act3D [206] takes a similar approach but build 3D feature field via sensed depth.
@@
Beyond context understanding, a few approaches explore the use of open-vocabulary properties of VLMs for state estimation [207, 199, 199, 208, 209]. Two such approaches, LEXIS [207] and FM-Loc [199], explore the use of CLIP [53] features to perform indoor localization and mapping. In particular, FM-Loc [199] leverages the vision-language grounding offered by CLIP and GPT-3 to detect objects and room labels of a query image, then uses that semantic information to match it with reference images. Similarly, LEXIS [207] builds a real-time topological SLAM graph where CLIP features are associated with graph nodes, enabling room-level scene recognition. Although these approaches display the potential of vision-language features for indoor place recognition, they do not explore the broad applicability of foundation model features. In this context, AnyLoc [208] explored the properties of dense foundation model features and combined them with unsupervised feature-aggregation techniques to achieve state-of-the-art place recognition, by large margin—anywhere, anytime, and under any view—showcasing broad applicability of self-supervised foundation model features for SLAM.
@@
Several works consider the notion of enabling robots to leverage the process of interactive perception, for extrapolating implicit knowledge about object properties in order to obtain performance improvements on downstream interactive robot learning tasks [210, 211, 212, 213, 214, 215, 167, 166, 200]. This process of interactive perception is often modeled after the way in which human infants first learn about the physical world—i.e., through interaction, and by learning representations of object concepts, such as weight and hardness, from the sensory information (haptic, auditory, visual) that is generated from those physical exploratory actions (e.g., grasping, lifting, dropping, pushing) on objects with diverse properties. In particular, MOSAIC [200] leverages LMMs to expedite the acquisition of unified multi-sensory object property representations; the authors show competitive performances of their framework in category recognition and ambiguous target-object fetch tasks, despite the presence of distractor objects, under zero-shot transfer conditions.
@@ 4.1.2 LLMs and VLMs in Task Planning
(a) Task plan generation via LLM and
grounding of visual states, SayCan [24]
(b) Task planning via code generation
LLM: Code as policy [216]
(c) VLMs for task planning and heuristic function,
plus a video dynamics model: VLP [217]
Figure 7: Examples of LLMs and VLMs used in task-level planning
@@
Task-level planning is to divide a complicated task into small actionable steps. In this case, we mainly talk about the agent planning on its own, in contrast to, e.g., using LLMs as a parser like Vision Language Navigation [218]. The agent needs to take intelligent sub-steps to reach the goal by interacting with the environment. SayCan [24] is a representative example of task-level planning: it uses LLMs to plan for a high-level task, e.g., “I spilled my drink, can you help?”. Then it gives concrete task plans like going to the counter, finding a sponge, and so on. Similarly, VLP [217] aims to improve the long-horizon planning approach with an additional text-to-video dynamics model. These task-level planning methods do not have to worry about the precise execution of those sub-tasks in the environment, since they have the luxury of utilizing a set of pre-defined / pre-trained skills, then using LLMs to simply find proper ways to compose skills to achieve desired goals. There are more papers in this category, for example: LM-ZSP [219] introduce this task-level granularity as actionable steps; Text2Motion [220] uses similar ideas and augments the success rate of language based manipulation task. Previous methods typically generate task plans in the form of text. Some works like ProgPrompt [221], Code as Policy [216], GenSim [222], etc. obtain task plans in the form of code generation using LLMs. Using code as a high-level plan has the benefit of expressing functions or feedback loops that process perception outputs and parameterize control primitive APIs. In addition, it can describe the spatial position of an object accurately. This improved compositionality saves time in collecting more primitive skills. It also prescribes precise values (e.g., velocities) to ambiguous descriptions like ’faster’ and ’to the left’, depending on the context. Therefore, due to these benefits, code appears to be a more efficient and effective task-level planning language than natural language. Other form of planning techniques such as express the high-level plans in Planning Domain Definition Language (PDDL)[223] also showed significant improvement on LLMs planning power over long horizon tasks, more on this will be discussed in the Section 4.1.6.
@@
In addition to use LLMs to direct generate plans, they are also used in searching and evaluating with external memory structure such as scene graph. SayPlan [224] employs 3D scene graph (3DSG) representations to manage the complexity of expansive environments. By exploiting hierarchical 3DSGs, LLMs can semantically search for relevant sub-graphs in multi-floor household environments, reducing the planning horizon and integrating classical path-planning to refine initial plans, iteratively. Reasoned Explorer [225] employs LLMs as evaluators to score each node in a 2D undirected graph. It uses this graph as a map to store both visited points and the frontiers’ LLM evaluations. These external memories and incremental map-building approach breaks the context length limit of using LLMs to generate long plans, which scales LLM-based navigation agents to large environments. One thing to note is that, although task-level planning is agnostic to physical embodiment, it does require grounding to a specific robot’s physical form (or “morphology”) and environment when deployed; grounding techniques will be covered in Section 4.1.4.

### v3

@@ 3 Challenges on General-purpose Robots
In this section, we summarize 5 core challenges that general-purpose robotic system face, each detailed in the following subsections. Whereas similar challenges have been discussed in prior literature (Section 1.2), this section mainly focuses on the challenges that may potentially be solved by appropriately leveraging foundation models, given the evidence from current research results. We also depict the taxonomy in the section for easier review in Figure 3.
Challenges in Robotics (3)Generalization (3.1)perceptionmorphologyData Scarcity (3.2)simulationreal-worldRequirements of Models (3.3)model-basedmodel-freeTask Specification (3.4) goal imagereward generationlanguage promptUncertainty and Safety (3.5) epistemicaleatoricprovable safety
@@ 3.1 Generalization
Robotics systems often struggle with accurate perception and understanding of their environment. Limitations in computer vision, object recognition, and semantic understanding made it difficult for robots to effectively interact with their surroundings. Traditional robotics systems often relied on analytic hand-crafted algorithms, making it challenging to adapt to new or unseen situations. They also lacked the ability to generalize their training from one task to another, further limiting their usefulness in real-world applications. This generalization ability is also reflected in terms of the generalization of planning and control in different tasks, environments, and robot morphologies. For example, specific hyperparameters for, e.g., classical motion planners and controllers need to be tuned for specific environments [49, 50, 51]; RL-based controllers are difficult to transfer across different tasks and environments [52, 53, 54]. In addition, due to differences in robotic hardware, it is also challenging to transfer models across different robot morphologies [55, 56]. By applying foundation models in robotics, the generalization problem is partially resolved, which will be discussed in the next Section 4. Further challenges, such as generalization across different robotic morphologies, remain demanding.
@@ 3.2 Data Scarcity
Data has always been the cornerstone of learning-based robotics methods. The need for large-scale, high-quality data is essential to develop reliable robotic models. Several endeavors have been attempted to collect large-scale datasets in the real world, including autonomous driving [1, 2, 57], robot manipulation trajectories [58, 59, 60], etc. Collecting robot data from human demonstration is expensive [29]. The diverse range of tasks and environments where robots are used even complicates the process of collecting adequate and extensive data in the real world. Moreover, gathering data in real-world settings can be problematic due to safety concerns [49]. To overcome these challenges, many works [61, 62, 63, 64, 65, 66] attempt generating synthetic data in simulated environments. These simulations offer realistic virtual worlds where robots can learn and apply their skills to real-life scenarios. Simulations also allow for domain randomization, as well as the potential to update the parameters of the simulator to better match the real world physics [49], helping robots to develop versatile policies. However, these simulated environments still have their limits, particularly in the diversity of objects, making it difficult to apply the learned skills directly to real-world situations. Collecting real-world robotic data with a scale comparable to the internet-scale image/text data used to train foundation models is especially challenging. One promising approach is collaborative data collection across different laboratories and robot types [67, 68, 69], as shown in Figure 4. However, an in-depth analysis of the Open-X Embodiment Dataset reveals certain limitations regarding data type availability. First, in Figure 4a, the robot morphology utilized for data collection is restrictive; out of the 73 datasets, 55 are dedicated to single-arm manipulation tasks. Only one dataset pertains to quadruped locomotion, and a single dataset addresses bi-manual tasks. Second, in Figure 4b, the predominant scene type for these manipulation tasks is tabletop setups, often employing toy kitchen objects. Such objects come with inherent assumptions including rigidity and negligible weight, which may not accurately represent a wider range of real-world scenarios. Third, in Figure 4c, our examination of data collection methods indicates a predominance of human expert involvement, predominantly through virtual reality (VR) or haptic devices. This reliance on human expertise highlights the challenges in acquiring high-quality data and suggests that significant human supervision is required. For instance, the RT-1 Robot Action dataset necessitated a collection period of 17 months, underscoring the extensive effort and time commitment needed for data accumulation with human involvement.
@@
Figure 4: Comprehensive visualizations of the Open-X Embodiment and Droid Dataset encompassing robot morphologies, environment types, and data collection methods.
@@ 3.3 Requirements of Models and Primitives
Classical planning and control methods usually require carefully engineered models of the environment and the robot. Optimal control methods require good dynamics models (i.e., world transition models) [8, 70]; motion planning requires a map of the environment [71], the states of the objects robots interact with [72], or a set of pre-defined motion primitives [73]; task planning requires pre-computed object classes and pre-defined rules [74, 75], etc. Previous learning-based methods (e.g., imitation and reinforcement learning) train policies in an end-to-end manner that directly gets control outputs from sensory inputs [58], avoiding building and using models. These methods partially solve this problem of relying on explicit models, but they often struggle to generalize across different environments and tasks. This raises two problems: (1) How can we learn model-agnostic policies that can generalize well? Or, alternatively, (2) How can we learn good world models so that we can apply classical model-based approaches? We see some recent works that aim to resolve these problems using foundation models (especially in a model-free manner), which will be systematically discussed in Section 4. However, the call for world models for robotics remains an intriguing frontier, which will be discussed in Section 6.
@@ 3.4 Task Specifications
Understanding the task specification and grounding it in the robot’s current understanding of the world is a critical challenge to obtaining generalist agents. Often, these task specifications are provided by users with limited understanding of the limitations on the robot’s cognitive and physical capabilities. This not only raises questions about what the best practices are for providing these task specifications, but also about the naturalness and ease of crafting these specifications. Understanding and resolving ambiguity in task specifications, conditioned on the robot’s understanding of its own capabilities, is also challenging. Foundation models are a promising solution for this challenge: task specification can be formulated as language prompts [29, 30, 26, 76], goal images [77], rewards for policy learning [28, 78], etc.
@@ 4 Review of Current Research Methodologies
In this section, we summarize the current research methodologies of foundation models for robotics.
Foundation Models for Robotics (4) Robotic Foundation Models (4.2)General-purpose (4.2.2) Single-purpose Action Generation Models (4.2.1)Vision and Language Pre-training Reinforcement Learning at Scale (Languange-Conditioned) Imitation LearningMotion Planning Models (4.2.1)(Vision Languange) Foundation Models used in Robotics (4.1)Prompting (4.1.6) Data Generation (4.1.5) Action Grounding (4.1.4) Robotic ModulesLLM and VLM in Action Generation (4.1.3) LLM and VLM in Task Planning (4.1.2) VFM and VLM in Perception (4.1.1)
In Section 4.1, we mainly discuss foundation models for robotics in two categories: Foundation Models used in Robotics and Robotics Foundation Models (RFMs). For Foundation Models used in Robotics, we mainly highlight applications of vision and language foundation models used in a zero-shot manner, meaning no additional fine-tuning or training is conducted. In Section 4.2, however, we mainly focus on Robotics Foundation Models, wherein these approaches may warm-start models with vision-language pre-trained initialization and/or directly train the models on robotics datasets. Figure 5 shows the detailed taxonomy of this section.
We review the methods presented in these papers following the convention of a typical robotic system which consists of perception, planning, and control modules. Here, we combine motion planning and control into one piece—action generation and treat motion planning modules as higher-level and control as lower-level action generation. It is important to notice that although most of the works use foundation models in different functionality modules of the robotic systems, we will classify these papers into categories based on the module to which the paper contributes the most. There are, however, certain applications of the vision and language foundation models that go across these robotics modules, e.g., grounding of these models in robotics, and generating data from LLMs and VLMs. Given the autoregressive nature of current LLMs, they often grapple with extended horizon tasks. Thus, we also delve into advanced prompting methods proposed in the literature to ameliorate this limitation and enhance planning power. We list these applications in sections 4.1.4, 4.1.5 and 4.1.6, as a different perspective to analyze these works.
@@
Figure 6: Here we plot some representative works of foundation models used in robotics, and robotic foundation models. The horizontal axis represents the complexity of input data; the vertical axis represents the complexity of output action space. The complexity of the input data is brought by the modality of the data, e.g. image data is more complex than the text data. The complexity of output action space is mainly determined by the output dimension of the foundation model in the robotic tasks.
@@ 4.1 Foundation Models used in Robotics
. In this section, we focus on zero-shot applications of vision and language foundation models in robotic tasks, which include mainly zero-shot deployment of VLMs used in robotic perception, in context learning of LLMs for task-level and motion-level planning, as well as action-generation.
@@ 4.1.1 VFMs and VLMs in Robot Perception
. Recently, the grounding of vision and language foundation models with geometric and object-centric representations of the world has enabled tremendous progress in context understanding, which is a vital requirement for robots to interact with the real world. We will thoroughly examine the application of VFMs and VLMs in robotic perception from various perspectives.
@@
Most of the previous works discussed above utilize only 2D representation of the objects and environment. To enrich the representation of foundation models in 3D space, F3RM [98], GNFactor [99] and GeFF [102] distill 2D foundation model features into 3D space, by combining with (generalizable) NeRF. In addition, GNFactor [99] also apply these distilled features in policy learning. Act3D [103] takes a similar approach but build 3D feature field via sensed depth.
@@
Beyond context understanding, a few approaches explore the use of open-vocabulary properties of VLMs for state estimation [104, 105, 105, 106, 107]. Two such approaches, LEXIS [104] and FM-Loc [105], explore the use of CLIP [23] features to perform indoor localization and mapping. In particular, FM-Loc [105] leverages the vision-language grounding offered by CLIP and GPT-3 to detect objects and room labels of a query image, then uses that semantic information to match it with reference images. Similarly, LEXIS [104] builds a real-time topological SLAM graph where CLIP features are associated with graph nodes, enabling room-level scene recognition. Although these approaches display the potential of vision-language features for indoor place recognition, they do not explore the broad applicability of foundation model features. In this context, AnyLoc [106] explored the properties of dense foundation model features and combined them with unsupervised feature-aggregation techniques to achieve state-of-the-art place recognition, by large margin—anywhere, anytime, and under any view—showcasing broad applicability of self-supervised foundation model features for SLAM. Extending this further, FoundLoc [107] coupled AnyLoc with a VIO pipeline to perform visual GNSS-denied localization. For the first time, this work showcased that VLMs can be deployed on resource-constrained unmanned aerial vehicles (UAVs) & embedded Jetson hardware for state estimation in the wild.
@@
Several works consider the notion of enabling robots to leverage the process of interactive perception [108], for extrapolating implicit knowledge about object properties in order to obtain performance improvements on downstream interactive robot learning tasks [109, 110, 111, 112, 113, 114, 56, 55, 115, 116]. This process of interactive perception is often modeled after the way in which human infants first learn about the physical world—i.e., through interaction, and by learning representations of object concepts, such as weight and hardness, from the sensory information (haptic, auditory, visual) that is generated from those physical exploratory actions (e.g., grasping, lifting, dropping, pushing) on objects with diverse properties. In particular, MOSAIC [115] leverages LMMs to expedite the acquisition of unified multi-sensory object property representations; the authors show competitive performances of their framework in category recognition and ambiguous target-object fetch tasks, despite the presence of distractor objects, under zero-shot transfer conditions. UncOS [116] queries LVMs to generate a distribution of object segmentation hypotheses. They convert the hypotheses into a distribution over world models and use that to select robot perturbation actions for embodied disambiguation of object segmentation.
@@ 4.1.2 LLMs and VLMs in Task Planning
Task-level planning is to divide a complicated task into small actionable steps. In this case, we mainly talk about the agent planning on its own, in contrast to, e.g., using LLMs as a parser like Vision Language Navigation [117]. The agent needs to take intelligent sub-steps to reach the goal by interacting with the environment. SayCan [26] is a representative example of task-level planning: it uses LLMs to plan for a high-level task, e.g., “I spilled my drink, can you help?”. Then it gives concrete task plans like going to the counter, finding a sponge, and so on. Similarly, VLP [118] aims to improve the long-horizon planning approach with an additional text-to-video dynamics model. These task-level planning methods do not have to worry about the precise execution of those sub-tasks in the environment, since they have the luxury of utilizing a set of pre-defined / pre-trained skills, then using LLMs to simply find proper ways to compose skills to achieve desired goals. There are more papers in this category, for example: LM-ZSP [119] introduce this task-level granularity as actionable steps; Text2Motion [120] uses similar ideas and augments the success rate of language based manipulation task. Previous methods typically generate task plans in the form of text, Some works like ProgPrompt [121], Code as Policy [122], GenSim [123], etc. obtain task plans in the form of code generation using LLMs. [124] take a different approach by obtaining relational constraints from code writing LLMs. Using code as a high-level plan has the benefit of expressing functions or feedback loops that process perception outputs and parameterize control primitive APIs. In addition, it can describe the spatial position of an object accurately. This improved compositionality saves time in collecting more primitive skills. It also prescribes precise values (e.g., velocities) to ambiguous descriptions like ’faster’ and ’to the left’, depending on the context. Therefore, due to these benefits, code appears to be a more efficient and effective task-level planning language than natural language. Other forms of planning techniques such as expressing the high-level plans in Planning Domain Definition Language (PDDL)[125] also showed significant improvement in LLMs planning power over long horizon tasks, more on this will be discussed in Section 4.1.6.
@@
In addition to using LLMs to directly generate plans, they are also used in searching and evaluating with external memory structures such as scene graph. SayPlan [126] employs 3D scene graph (3DSG) representations to manage the complexity of expansive environments. By exploiting hierarchical 3DSGs, LLMs can semantically search for relevant sub-graphs in multi-floor household environments, reducing the planning horizon and integrating classical path-planning to refine initial plans, iteratively. Reasoned Explorer [127] employs LLMs as evaluators to score each node in a 2D undirected graph. It uses this graph as a map to store both visited points and the frontiers’ LLM evaluations. These external memories and incremental map-building approaches break the context length limit of using LLMs to generate long plans, which scales LLM-based navigation agents to large environments. One thing to note is that, although task-level planning is agnostic to physical embodiment, it does require grounding to a specific robot’s physical form (or “morphology”) and environment when deployed; grounding techniques will be covered in Section 4.1.4.